What is the pressure of 1.75 moles of a gas that occupies 35 L at -50 ˚C?


1
Expert's answer
2020-11-09T13:53:03-0500

Q142580

What is the pressure of 1.75 moles of a gas that occupies 35 L at -50 ˚C?

Solution:

In question, we are given,

moles of gas, n = 1.75moles

The Volume occupied by gas, V = 35L

Temperature of gas, T = -500C.

Gas constant, R = 0.08206 L-atm/mol-K



We can use this information in the ideal gas formula, PV = nRT, and find the Pressure of the gas.

In the ideal gas equation, we must use Temperature in kelvin.

Convert -500C to Kelvin by using the formula,


"T (K) = T(^0C ) + 273.15"


T(K) = -50 + 273.15 = 223.15K


Substitute all the given information in the ideal gas equation, we have



"PV = nRT"

P * 35L = 1.75mol * 0.08206 L-atm/mol-K * 223.15K ;


P * 35L = 32.045 L-atm

divide both the side by 35L we have


P*35L/35L = 32.045 L-atm/35L


P = 0.9156 atm ;


In question, the quantity with the least number of significant figures is 35L. It is in 2 significant figures.

So our final answer must also be in 2 significant figures.


0.9156 atm in 2 significant figures can be written as 0.92 atm.


Hence, the pressure of the gas is 0.92 atm.